### v1

# Rational torsion on hyperelliptic jacobian varieties

Abstract.
Fixing an integer $g\geq 2$, it was conjectured by Flynn that there exists a constant $\kappa$, independent of $g$, such that for every $m\leq\kappa g$, there exists a hyperelliptic curve of genus $g$ over ${\mathbb{Q}}$ with a rational $m$-torsion point on its Jacobian. Leprévost proved this conjecture with $\kappa=3$. In this work we prove that given an integer $N$ in the interval $[3g,4g+1]$, $g\geq 3$, satisfying certain partition conditions, there exist infinite families of hyperelliptic Jacobian varieties with a rational torsion point of order $N$. In particular, we establish the existence of such varieties for $N=4g+1$ when $g$ is odd and for $N=4g-1$ when $g$ is even. A few explicit applications of this result produce the first known infinite examples of torsion $13$ when $g=3$, torsion $15$ when $g=4$, and torsion $17,18,21$ when $g=5$. In fact, we show that infinitely many of the latter abelian varieties are absolutely simple.
footnotetext: Keywords: Hyperelliptic curves, Jacobian varieties, torsion
2020 Mathematics Subject Classification: 11G30, 14H25
1. Introduction
Given an integer $n\geq 1$, the existence of a uniform bound, that depends solely on $n$, on the size of the torsion subgroups of elliptic curves defined over a number field of degree $n$ is guaranteed by a result of Merel, [22]. Complete lists for the structure of these torsion subgroups are explicitly given when $n=1,2,3$ in [21, 10, 11, 6]. For abelian varieties of fixed dimension $d\geq 2$, not only such a uniform boundedness result is far from reach, but no complete classification of possible torsion orders has been realized over any number field.
A computational approach toward studying torsion subgroups of abelian varieties of dimension $d\geq 2$ is to find integers that are realized as orders of torsion points on these varieties. For example, a family of genus $2$ curves whose Jacobian varieties possess torsion points of order $\ell$ over ${\mathbb{Q}}$ was given in [7] when $\ell=11$, and in [13, 14, 16], when $\ell=13$, $15$, $17$, $19$ or $21$. In [9], several explicit families of curves of genus $2$ and $3$ whose Jacobians possess large rational torsion subgroups were displayed.
Researchers questioned the possibility of constructing families of abelian varieties in which the size of the rational torsion subgroup of the variety increases as the dimension of the variety increases. In [8], fixing an integer $g\geq 2$, it was proved that any integer in the interval $[g^{2}+2g+1,g^{2}+3g+1]$ is realized as the order of a ${\mathbb{Q}}$-rational torsion point on the Jacobian of a hyperelliptic curve of genus $g$. Other families of hyperelliptic Jacobian varieties of dimension $g$ with ${\mathbb{Q}}$-rational torsion points of quadratic order in $g$ were given in [12, 15, 18]. In [7], it was conjectured that there exists a constant $\kappa$, independent of $g$, such that for every $m\leq\kappa g$, there exists an infinite family of hyperelliptic curves of genus $g$ over ${\mathbb{Q}}$ with a torsion point of order $m$ on its Jacobian. Leprévost proved this conjecture with $\kappa=3$ in [16].
Since $11$ is the smallest prime for which there is no elliptic curve over ${\mathbb{Q}}$ with rational point of order the prime, finding abelian varities with torsion order $11$ was an interesting challenge for researchers. Jacobians of genus $2$ curves with torsion order $11$ over ${\mathbb{Q}}$ can be found in [7, 2, 4, 5]; Flynn’s conjecture, later proved by Leprévost, establishes the existence of infinitely many such varieties when $g\geq 4$. Recently, Jacobians of genus $3$ hyperelliptic curves with a ${\mathbb{Q}}$-rational torsion point of order $11$ was given in [4, 6]. This makes $13$ the next challenging prime for abelian varieties of dimension $g$ when $g=3,4$.
In this work, we extend the interval given by Leprévost to prove Flynn’s conjecture to include new integers that are realized as torsion orders for hyperelliptic Jacobian varieties. The divisor at infinity has finite order in the Jacobian of a hyperelliptic curve if and only if a certain continued fraction expansion of a power series is periodic. We show that for every integer $N$ in the interval $[3g,4g+1]$, $g\geq 3$, satisfying certain partition conditions, there exist infinite families of hyperelliptic curves of genus $g$ whose Jacobian varieties have a ${\mathbb{Q}}$-rational torsion point of order $N$. These partition conditions appear due to the fact that the order of the divisor at infinity in the Jacobian is the sum of the degrees of some partial quotients in the corresponding continued fraction expansion. Some of the integers satisfying these partition conditions are $14g+$ 1 when $g$ is odd; and $14g-$ 1 when $g$ is even. Consequently, we produce infinite families of hyperelliptic Jacobian varieties with torsion order $N=13$ when $g=3$, torsion $15$ when $g=4$, and torsion $17,18,21$ when $g=5$. These are the first examples of infinitely many varieties corresponding to the aforementioned pairs $(g,N)$. For a survey of all torsion points of a certain order that have been found for Jacobian varieties of hyperelliptic curves of genus $2,3$, and $4$, the reader may consult [23]. In addition, we show that infinitely many of the Jacobian varieties that we construct are absolutely simple.
Acknowledgment
All the calculations in this work were performed using Magma, [3], and Mathematica, [25]. M. Sadek is supported by The Scientific and Technological Research Council of Turkey, TÜBİTAK, research grant ARDEB 1001/122F312.
2. Torsion divisors at infinity
Throughout this section $k$ is a field with $\operatorname{char}k\neq 2$. We let $\overline{k}$ be an algebraic closure of $k$.
2.1. Continued fractions in $K[x]$
Let $K=k(x)$ be function field over the field $k$. Define a function
$\varphi:K-k[x]\rightarrow K,\qquad\varphi(\alpha)=1/(\alpha-[\alpha])$
where $[\alpha]$ is the polynomial part of $\alpha$. For positive integers $r$ let $\varphi_{r}$ be the $r$-fold composition of $\varphi$ with itself and let $\varphi_{0}$ be the identity map, namely,
$\alpha_{0}=\alpha,\quad\alpha_{r}=\varphi_{r}(\alpha),\quad\alpha_{r+1}=1/\left(\alpha_{r}-\left[\alpha_{r}\right]\right)\quad\text{ for }r\geqslant 1.$
In addition, $a_{r}:=\left[\alpha_{r}\right]$ is called the partial quotient of $\alpha$. Now the continued fraction expansion of $\alpha$ is given by
$\alpha=a_{0}+\cfrac{1}{a_{1}+\cfrac{1}{a_{2}+{}_{\ddots}}}.$
One can write the latter continued fraction in short as $\alpha=\left[a_{0},a_{1},a_{2},\ldots\right]$.
Definition 2.1.
An even length sequence $(a_{1},\cdots,a_{2r})$ is said to be skew symmetric of skew value $\gamma\in k-\{0\}$ if $a_{2(r-i)}=c_{i}a_{2i+1}$, $0\leq i<r$, where $c_{i}=\gamma$ if $i$ is even; and $c_{i}=1/\gamma$ otherwise.
2.2. The divisor at infinity
Let $C$ be a hyperelliptic curve of genus $g\geq 2$ over $k$ given by the equation $y^{2}=f(x)$ where $f(x)\in k[x]$ is of even degree $22g+$ 2, that has no repeated root in $\overline{k}$ and the leading coefficient of $f(x)$ is a nonzero square in $k$. Let $\infty_{+}$ and $\infty_{-}$ be the two $k$-rational points at infinity on $C$.
There exists a natural embedding of the curve $C$ into its Jacobian $J$ that maps a point $P$ to the divisor class $[P-D]$, where $D$ is a fixed divisor of degree $1$. This map restricts to $C(k)\xhookrightarrow{}J(k)$ if $D$ is a $k$-rational divisor. In particular, if there exists a point $P_{0}\in C(k)$, one can choose $D$ to be the rational divisor $P_{0}$.
The divisor at infinity on $C$ is the divisor $D_{\infty}=\infty_{+}-\infty_{-}$. It defines a point in $J(k)$. We say that $D_{\infty}$ is torsion of order $N$ if its class in $J$ has order $N$. The following result is due to Adams and Razar, see [1, Theorem 5.1].
Proposition 2.2.
Let $C$ be a hyperelliptic curve described as above. The divisor $D_{\infty}$ is torsion if and only if the continued fraction of $y$ is periodic.
In the latter case, the following statements hold:
  • a)
    $y=[a_{0};\overline{a_{1},a_{2},\cdots,a_{m-1},2\gamma a_{0},a_{m-1},\cdots,a_{2},a_{1},2a_{0}}]$, where $(a_{1},a_{2},\cdots,a_{m-1})$ is skew symmetric of skew value $\gamma^{-1}$.
  • b)
    $\deg a_{0}=g+1$, and $1\leq\deg a_{i}\leq g$ for $11\leq i\leq m-$ 1.
  • c)
    The order of $D_{\infty}$ is $g+1+\sum_{i=1}^{m-1}\deg a_{i}$.
The properties in Proposition 2.2 can be found in [1] and [24, Theorem 4.2].
Remark 2.3.
In Proposition 2.2, we call (the minimal) $m$ the quasi-period length. The (minimal) period length $n$ is equal to the quasi-period length $m$ exactly when $\gamma=1$. Further, if $m\neq n$, then $n=2m$ and $m$ must be odd. In the latter case, $m$ is called a strict quasi-period length.
3. Construction of curves
Fixing an integer $m\geq 0$, we set ${\mathbb{Q}}(\textbf{w}):={\mathbb{Q}}(w_{1},\cdots,w_{m})$, where $w_{1},\cdots,w_{m}$ are algebraically independent. In this section, we construct hyperelliptic curves $C$ of genus $g$ over ${\mathbb{Q}}(\textbf{w})$ described by the equation
$y^{2}=h_{2g+2}(\textbf{w})^{2}x^{2g+2}+h_{2g+1}(\textbf{w})x^{2g+1}+\cdots+h_{0}(\textbf{w}),\qquad h_{i}(\textbf{w})\in{\mathbb{Q}}(\textbf{w}),\quad i=0,\cdots,2g+2.$
The curve $C$ admits two ${\mathbb{Q}}(\textbf{w})$-rational points at infinity denoted by $\infty^{+}$ and $\infty^{-}$.
We investigate the case when the divisor $D_{\infty}=\infty^{+}-\infty^{-}$ in the Jacobian of $C$ is torsion with the quasi-period length $m$ is the same as period length $n$ and both are equal to $6$. This means that the continued fraction expansion of $y$ is given by
(1) $[a_{0}(x);\overline{a_{1}(x),a_{2}(x),a_{3}(x),a_{2}(x),a_{1}(x),2a_{0}(x)}]$
where $\deg a_{0}=g+1$ and $1\leq\deg a_{i}\leq g$ when $i=1,2,3$.
In order to construct our desired curves, a solution of a certain polynomial Diophantine equation must be found.
Lemma 3.1.
Let $r(x),q(x),a_{1}(x),a_{2}(x),a_{3}(x)\in{\mathbb{Q}}[x]$. The polynomial Diophantine equation
$-2q(x)-2a_{2}(x)+a_{1}(x)r(x)+a_{3}(x)r(x)-2q(x)a_{1}(x)a_{2}(x)+a_{1}(x)a_{2}(x)a_{3}(x)r(x)-a_{2}(x)^{2}a_{3}(x)\equiv 0$
has the following solution
$q(x)$ $=$ $u(x)r(x)/2,$
$a_{2}(x)$ $=$ $a_{1}(x)r(x),$
$a_{3}(x)$ $=$ $a_{1}(x)^{2}u(x)r(x)+a_{1}(x)+u(x),$
where $u(x)\in{\mathbb{Q}}[x]$.
Proof: One may check that the claimed solution satisfies the polynomial Diophantine equation. $\Box$
Theorem 3.2.
Fix an integer $g\geq 3$. Let $\alpha,\beta\geq 1,\gamma\geq 0$ be integers such that $12\alpha+2\beta+\gamma=g+$ 1. Let $a_{1}(x),r(x),u(x)\in{\mathbb{Q}}[x]$ be of degrees $\alpha,\beta,\gamma$, respectively. If the affine equation
$y^{2}=r(x)^{2}\left(u(x)(a_{1}(x)^{2}r(x)+1)+a_{1}(x)\right)^{2}+4\left(u(x)a_{1}(x)r^{2}(x)+r(x)\right)$
describes a hyperelliptic curve $C$, then the divisor at infinity is torsion of order $g+1+6\alpha+3\beta+\gamma$.
In addition, if $\beta=1$, and the Galois group of the polynomial
$r(x)\left(u(x)(a_{1}(x)^{2}r(x)+1)+a_{1}(x)\right)^{2}+4\left(u(x)a_{1}(x)r(x)+1\right)$
is either the full symmetric group $\operatorname{\textbf{S}}_{2g+1}$ or the alternating group $\operatorname{\textbf{A}}_{2g+1}$, then the Jacobian $J_{C}$ of $C$ satisfies that $\operatorname{End}(J_{C})={\mathbb{Z}}$, in particular, $J_{C}$ is an absolutely simple abelian variety
Proof: In view of (1), a hyperelliptic curve $C$ for which the divisor at infinity is torsion with $m=n=6$ is described by an equation of the form
$y^{2}=a_{0}^{2}+\frac{a_{2}(2+a_{2}a_{3})+2a_{0}(1+a_{2}a_{3}+a_{1}a_{2}(2+a_{2}a_{3}))}{(1+a_{1}a_{2})(a_{3}+a_{1}(2+a_{2}a_{3}))}.$
In order for the latter equation to be a polynomial equation, the latter fraction must be a polynomial $h\in{\mathbb{Q}}[x]$. This yields that
$a_{0}=\frac{a_{1}h}{2}+\frac{-a_{2}^{2}a_{3}+(a_{1}+a_{3})h+a_{2}(-2+a_{1}a_{3}h)}{2(1+a_{2}a_{3}+a_{1}a_{2}(2+a_{2}a_{3}))}.$
By comparing the degrees of the numerator and denominator of the latter fraction as polynomials in $a_{1},a_{2}$ and $a_{3}$, we may assume that $h=2qa_{2}+r$ with $\deg r<\deg q$ to obtain that
$a_{0}=\frac{a_{1}h+2q}{2}+\frac{-2a_{2}-2q+a_{1}r+a_{3}r-2qa_{1}a_{2}+a_{1}a_{2}a_{3}r-a_{2}^{2}a_{3}}{2(1+a_{2}a_{3}+a_{1}a_{2}(2+a_{2}a_{3}))}.$
We notice that if the degree of $q(x)$ is strictly less than the sum of the degrees of $a_{2}(x)$ and $a_{3}(x)$, then the numerator of the latter fraction is of strictly less degree than the denominator. Therefore, for $a_{0}$ to be a polynomial in ${\mathbb{Q}}[x]$, we may assume that $-2a_{2}-2q+a_{1}r+a_{3}r-2qa_{1}a_{2}+a_{1}a_{2}a_{3}r-a_{2}^{2}a_{3}\equiv 0\in{\mathbb{Q}}[x]$. According to Lemma 3.1, a solution to this polynomial Diophantine equation is provided giving rise to the affine equation $y^{2}=(a_{1}h+2q)^{2}/4+h$. After using the explicit description of the solution in Lemma 3.1, the curve $C$ is described by
$y^{2}=r(x)^{2}\left(u(x)(a_{1}(x)^{2}r(x)+1)+a_{1}(x)\right)^{2}+4\left(u(x)a_{1}(x)r^{2}(x)+r(x)\right).$
Now, when $\beta=1$, the rest of the statement will follow from [26, Theorem 1.3]. $\Box$
Remark 3.3.
In Theorem 3.2, we insist that $\beta:=\deg r\geq 1$, since otherwise $a_{3}=2a_{0}$. This means that in (1), the quasi period length, $m$, and the period length, $n$, satisfy that $m=n=3$.
We notice that in Theorem 3.2, if $a_{1}(x)=\sum_{i=0}^{\alpha}A_{i}x^{i}$, $r(x)=\sum_{i=1}^{\beta}R_{i}x^{i}$, and $u(x)=\sum_{i=0}^{\gamma}U_{i}x^{i}$, then the curve $C$ is defined by an equation of the form $y^{2}=f(x)$ where
$f(x)\in{\mathbb{Q}}[A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma}][x].$
The discriminant of $C$ is a constant multiple of the discriminant $\Delta_{f}$ of the polynomial $f$, see for example [20] or [19], hence it is a polynomial in ${\mathbb{Q}}[A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma}]$. If we consider the affine space $\mathbb{A}^{\alpha+\beta+\gamma+3}$ with coordinates $A_{0},\cdots A_{\alpha}$, $R_{0},\cdots R_{\beta}$, $U_{0},\cdots,U_{\gamma}$, then for a choice $(A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma})\in\mathbb{A}^{\alpha+\beta+\gamma+3}$, the equation $y^{2}=f(x)$ defines a hyperelliptic curve of genus $g$ over ${\mathbb{Q}}$ if and only if $(A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma})$ is chosen to lie in the complement of the hypersurface $\Delta_{f}=0$ in the affine space $\mathbb{A}^{\alpha+\beta+\gamma+3}$.
As a direct consequence of Theorem 3.2, we get the following result.
Corollary 3.4.
Fix an integer $g\geq 3$. Let $N$ be an integer such that $13g\leq N\leq 4g+$ 1. If there exists integers $\alpha,\beta\geq 1,\gamma\geq 0$ such that $12\alpha+2\beta+\gamma=g+$ 1 and $N=g+1+6\alpha+3\beta+\gamma$, then there exists infinitely many hyperelliptic curves of genus $g$ over ${\mathbb{Q}}$ whose Jacobian posses a rational point of order $N$. In particular, if $g$ is odd, respectively even, there are infinitely many hyperelliptic curves of genus $g$ over ${\mathbb{Q}}$ whose Jacobian posses a rational point of order $14g+$ 1, respectively $14g-$ 1.
Although our trials to study hyperelliptic curves $C$ for which the corresponding continued fraction in Proposition 2.2 has quasi-period length $m\leq 5$ yielded torsion divisors at infinity with order $\operatorname{O}(kg)$, where $k<3$, we successfully obtained families of hyperelliptic curves whose jacobain contains torsion points whose order is not listed in literature. As an example, when $m=n=4$, one obtains the following equation
$y^{2}=\left(rqa_{1}^{2}-q^{2}a_{1}+ra_{1}+q\right)^{2}+4(rqa_{1}-q^{2}+r).$
Choosing $r,q$ to be ${\mathbb{Q}}$-rationals and $q^{2}\neq r$ when $g$ is odd; and $q$ to be ${\mathbb{Q}}$-rational whereas $r$ to be a linear polynomial if $g$ is even, the order of the torsion divisor at infinity becomes $25(g+1)/$ 2 if $g$ is odd; and $25g/2+$ 2 if $g$ is even.
Example 3.5.
Consider the following family of hyperelliptic curves of genus $3$
$C_{t}:\,y^{2}=(2(x^{2}+t)^{2}+(x^{2}+t)+1)^{2}+4(2(x^{2}+t)+1)$
over ${\mathbb{Q}}(t)$. Since the discriminant of $C_{t}$ is given by $2^{48}\cdot 5^{4}\cdot(1+t)(5+5t+4t^{3})$, if $t\neq-1$, then the class of the divisor $-(1:-2:0)$ ( 1 : - 2 : 0 ) on the Jacobian of $C_{t}$ is of order $10$.
4. Explicit families
In this section we provide explicit parametric families of hyperelliptic curves of genus $g\geq 3$ whose Jacobian possess ${\mathbb{Q}}$-rational torsion points of orders that does not appear in the literature.
Theorem 4.1.
There exists infinitely many hyperelliptic curves of genus $3$ over ${\mathbb{Q}}$ whose Jacobians posses a rational point of order $13$.
Proof: In Theorem 3.2, pick $r(x)=ax+b$, $a\neq 0$, to be a linear polynomial in ${\mathbb{Q}}[x]$, $u(x)=u\in{\mathbb{Q}}\setminus\{0\}$, and $a_{1}(x)=cx+d\in{\mathbb{Q}}[x]$, $c\neq 0$. When the discriminant of this curve is nonzero, we get a $5$-parametric family of hyperelliptic curves satisfying the hypothesis of the theorem, where that divisor at infinity is $(1:a^{2}\cdot c^{2}\cdot u:0)-(1:-a^{2}\cdot c^{2}\cdot u:0)$. $\Box$
In a similar fashion, one may prove the following theorem.
Theorem 4.2.
There exists infinitely many hyperelliptic curves of genus $4$ over ${\mathbb{Q}}$ whose Jacobians posses a rational point of order $15$.
Example 4.3.
If $u\neq 0$, then the hyperelliptic curve $C^{13}_{u,t}$ of genus $3$ described by the equation
$y^{2}=(x+t)^{2}(2t+u+x+u(x+t)(x+2t)^{2})^{2}+4(t+x+u(x+t)^{2}(x+2t))$
satisfies that the divisor at infinity is torsion of order $13$.
Example 4.4.
Consider the genus $4$ hyperelliptic curve described by
$C^{15}_{s,t}:\,y^{2}=(x+t)^{2}\left(x+(x-s)(1+x^{2}(x+t))\right)^{2}+4\left(t+x+x(-s+x)(t+x)^{2}\right).$
For a choice of $s,t$ such that the discriminant of $C_{s,t}$ is nonzero, one sees that the divisor at infinity $-(1:-1:0)$ ( 1 : - 1 : 0 ) is torsion of order $15$.
Remark 4.5.
In examples 4.3, the equation describing the curve $C^{13}_{u,t}$ is of the form $(x+t)f_{u,t}(x)$ where the polynomial $f_{u,t}$ has generic Galois group $\textbf{S}_{7}$. The latter statement can be justified by seeing that the specialization $f_{1,1}$ has Galois group $\textbf{S}_{7}$. It follows that $f_{u,t}$ has Galois group $\textbf{S}_{7}$ for infinitely many pairs $u,t$. According to Theorem 3.2, this means that the Jacobian $J_{C^{13}_{u,t}}$ of $C^{13}_{u,t}$ satisfies that $\operatorname{End}(J_{C^{13}_{u,t}})={\mathbb{Z}}$ and that $J_{C^{13}_{u,t}}$ is an absolutely simple abelian variety for infinitely many values of $u,t$.
The same argument yields that the Jacobian of $C^{15}_{s,t}$ in Example 4.4 is an absolutely simple variety for infinitely many choices of $s,t$.
We notice that for $g\geq 5$, there are at least two integers $N$ in the interval $[3g,4g+1]$ when $g$ is odd, $[3g,4g-1]$ when $g$ is even, satisfying the hypotheses of Corollary 3.4.
Example 4.6.
Consider the following families of hyperelliptic curves of genus $5$
$C^{17}_{s,t}$ ::: $y^{2}=4(t+x+x(t+x)^{2}(x^{2}+s))+(x+t)^{2}(x+(x^{2}+s)(1+x^{2}(x+t)))^{2}$
$C^{18}_{s,t,u}$ ::: $y^{2}=(x^{2}+s)^{2}(t+u+x+u(x+t)^{2}(x^{2}+s))^{2}+4(s+x^{2}+u(x+t)(x^{2}+s)^{2})$
$C^{21}_{s,t,u}$ ::: $y^{2}=4(s+x+u(x+s)^{2}(x^{2}+t))+(x+s)^{2}(t+u+x^{2}+u(x+s)(x^{2}+t)^{2})^{2}$
corresponding to the triple $(\alpha,\beta,\gamma)=(1,1,2),(2,1,0),(1,2,0)$, respectively, in Theorem 3.2. In fact, for the triple $(1,1,2)$, we chose $a_{1}(x)=x$, $r(x)=x+t$, $u(x)=x^{2}+s$, which implies that the divisor at infinity is torsion of order $17$.
For the triple $(1,2,0)$, we chose $a_{1}(x)=x+t$, $r(x)=x^{2}+s$, $u(x)=u\in{\mathbb{Q}}\setminus\{0\}$, yielding that the divisor at infinity is torsion of order $18$. Finally, for the triple $(2,1,0)$, we chose $a_{1}(x)=x^{2}+t$, $r(x)=x+s$, $u(x)=u\in{\mathbb{Q}}\setminus\{0\}$, with the divisor at infinity being torsion of order $21$.
The same argument as in Remark 4.5 shows that the Jacobians of the curves $C_{s,t}^{17}$ and $C^{21}_{s,t,u}$ are absolutely simple varieties for infinitely many choices of $s,t,u$.
Conflict of Interest statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data availability statement
The authors declare that the data supporting the findings of this study are available within the paper.